OpenStudy (anonymous):

A fan club that comprises of boys and girls have a total of 160 members. If 32 more girls join the club, the percentage of girls will increase by 5%. Find the number of boys in the club.

OpenStudy (anonymous):

let suppose no. of girls be y and no. of boys be x. So, x+y=160 now let %age of girls be z then, y*100/160=z by adding 32 girls, (y+32)100/(160+32)=z+5 by solving this we get y=62 by this 62+x=160 x=98

OpenStudy (anonymous):

Wait... but somehow the answer says the no of boys is 48.:(

OpenStudy (anonymous):

sorry sorry sorry.................... by mistake i calculated it wrong...............sorry for that again......ans is 48

OpenStudy (anonymous):

Nvm... haha..Thank you :DD